# Cambridge Authors

### v1

@@ 1 Introduction
Reinforcement learning (RL) enables agents to learn environment interaction through trial and error, aiming to maximize cumulative reward or achieve predefined goals [712192, doi:10.1287/ijoc.1080.0305]. Through this, RL makes agents learn how to learn. In recent years, reinforcement learning (RL) has driven substantial progress across a wide range of domains, including game theory [Szita2012], robotic control [DBLP:journals/corr/abs-1812-03201, 7758092], autonomous driving [9351818, DBLP:journals/corr/Shalev-ShwartzS16a], recommendation systems [CHEN2023110335], and natural language processing [DBLP:journals/corr/HeCHGLDO15, FERREIRA2015256]. In RL-based robotic control frameworks [DBLP:journals/corr/abs-1812-00568], agents observe the environment and take actions based on their current policy.
@@
Proximal Policy Optimization (PPO) [DBLP:journals/corr/SchulmanWDRK17] is a widely used policy gradient method. It stabilizes training by clipping probability ratios, thereby approximately enforcing a trust-region constraint. By constraining policy updates, PPO avoids the instability common in standard policy gradient methods. PPO has been applied to robotics [DBLP:journals/corr/abs-2109-11978] and is now also used in multimodal agent interaction [11044689]. However, in quadrupedal locomotion control, standard PPO often exhibits significant limitations. The learned policy often converges to suboptimal gaits, getting stuck in poor local optima [sutton2018reinforcement]. This issue is particularly pronounced in complex, uneven terrain, where learned policies often converge to suboptimal gaits and become trapped in poor local optima [sutton2018reinforcement]. For this, we improve a framework based on entropy.
@@
Incorporating an entropy bonus encourages exploration and helps prevent premature convergence [DBLP:journals/corr/abs-1801-01290]. Too much entropy induces excessive randomness in actions, which harms control accuracy and compromises system stability [ziebart2008maximum]. Conversely, too little entropy reduces behavioral diversity, leading to premature convergence and degraded policy performance. To balance exploration and exploitation, we propose dynamically adjusting the entropy coefficient during training. Through this our policy can adjust by itself. Motion smoothness also plays a critical role in quadrupedal locomotion, as legged systems are highly sensitive to abrupt changes in joint trajectories. Discontinuous actions can destabilize the robot’s center of mass, leading to imbalance or falls. We mitigate balance loss from discontinuous commands by constraining action changes across time steps, ensuring physically plausible motion. Smooth transitions between gaits allow the robot to adapt rapidly and stably to terrain changes or pushes. for this, robots can run smoothly in different environments. In many reinforcement learning problems, extrinsic rewards are sparse or delayed [DBLP:journals/corr/abs-1810-12894, DBLP:journals/corr/BellemareSOSSM16]. To address this, intrinsic rewards are commonly used to promote exploration.
@@ 3 Methods
3.1 Adaptive Entropy-Controlled Policy Optimization Mechanism
@@
3.2 Motion Continuity Regularization Framework
The motion of a robot is continuous. To ensure system stability its dynamic behavior must satisfy constraints. In our experiments, we observed that action sequences containing sharp transitions or high-frequency components frequently led to loss of balance and unstable locomotion.
@@ 3.2 Motion Continuity Regularization Framework
3.3 Intrinsic Motivation-Driven Exploration Enhancement Mechanism
@@ 4.1 Experimental Setup
We adopt the original PPO configuration of Rudin et al. for ANYmal in Isaac Gym as our baseline. For comparison, we use the benchmark source PPO with the same architecture, observation and action spaces. For controlled reasons, we compare the series of benchmarks based on the same pipeline, and our model differs only by the addition of adaptive entropy control, motion continuity regularization and intrinsic motivation components. Under this design, any systematic difference in performance between PPO and ECIM can be attributed to the proposed pipeline rather than differences in architecture, training budget or simulator settings.
@@ 4.2 Evaluation Metrics
4.3 Overall Performance of ECIM vs PPO
@@
4.3.1 Learning Efficiency and Task Performance
@@
4.3.2 Motion Stability and Smoothness
@@
4.3.3 Energy Efficiency
@@
4.4 Ablation Study of ECIM Components
@@
4.4.1 Overall Comparison
@@
4.4.2 Impact of AECPOM
@@
4.4.3 Impact of MCRF
@@
4.4.4 Impact of IMDEEM
@@
4.4.5 Attribution and Cross-Terrain Robustness
@@ 5 Conclusion
Author Contributions Wanru Gong and Xinyi Zheng conceived and designed the study. Wanru Gong and Xinyi Zheng conducted data gathering. GH performed statistical analyses. Wanru Gong, Xinyi Zheng, Xiaopeng Yang and Xiaoqing Zhu wrote the article.

### v2

@@ 1 Introduction
In recent years, mobile robots have garnered significant attention for their ability to autonomously carry out tasks across diverse environments [10.1115/1.4039394]. Among these, quadruped robots—equipped with bio-inspired legged architectures [10.1108/IR-11-2016-0310] or legged walking robots [CHEN2024105464]—exhibit remarkable agility and environmental adaptability, enabling stable and animal-like locomotion over complex and dynamic terrains [10552939, ms-9-1-2018]. Furthermore, compared to other legged robot configurations, quadrupeds strike an advantageous balance between structural stability and controllability, rendering them particularly well-suited for a wide range of real-world applications. Reinforcement learning (RL) enables agents to learn environment interaction through trial and error, aiming to maximize cumulative reward or achieve predefined goals [712192, doi:10.1287/ijoc.1080.0305]. Through this, RL makes agents learn how to learn. In recent years, reinforcement learning (RL) has driven substantial progress across a wide range of domains, including game theory [Szita2012], robotic control [DBLP:journals/corr/abs-1812-03201, 7758092], autonomous driving [9351818, DBLP:journals/corr/Shalev-ShwartzS16a], recommendation systems [CHEN2023110335], and natural language processing [DBLP:journals/corr/HeCHGLDO15, FERREIRA2015256]. In RL-based robotic control frameworks [DBLP:journals/corr/abs-1812-00568], agents observe the environment and take actions based on their current policy.
@@
Proximal Policy Optimization (PPO) [DBLP:journals/corr/SchulmanWDRK17] is a widely used policy gradient method. It stabilizes training by clipping probability ratios, thereby approximately enforcing a trust-region constraint. By constraining the size of each policy update, PPO helps reduce the instability often seen in standard policy gradient methods. It has been widely used in robotics tasks [DBLP:journals/corr/abs-2109-11978] and more recently in multi-model agent interaction [11044689]. In quadrupedal locomotion, however, several studies have shown that PPO still faces clear limitations, particularly in achieving stable gait transitions and consistent tracking performance [9765709, 4054408]. The learned policy often converges to suboptimal gaits, getting stuck in poor local optima [sutton2018reinforcement]. This issue is particularly pronounced in complex, uneven terrain, where learned policies often converge to suboptimal gaits and become trapped in poor local optima [sutton2018reinforcement]. For this, we improve a framework based on entropy.
@@
Incorporating an entropy bonus encourages exploration and helps prevent premature convergence [DBLP:journals/corr/abs-1801-01290]. Too much entropy induces excessive randomness in actions, which harms control accuracy and compromises system stability [ziebart2008maximum]. Conversely, too little entropy reduces behavioral diversity, leading to premature convergence and degraded policy performance. To balance exploration and exploitation, we propose dynamically adjusting the entropy coefficient during training. Through this our policy can adjust by itself. Motion smoothness also plays a critical role in quadrupedal locomotion, as legged systems are highly sensitive to abrupt changes in joint trajectories [10354849, Wang_Wang_Zhang_Dai_2022]. Discontinuous actions can destabilize the robot’s center of mass, leading to imbalance or falls [6696874, 6518581]. We mitigate balance loss from discontinuous commands by constraining action changes across time steps, ensuring physically plausible motion. Smooth transitions between gaits allow the robot to adapt rapidly and stably to terrain changes or pushes. for this, robots can run smoothly in different environments. In many reinforcement learning problems, extrinsic rewards are sparse or delayed [DBLP:journals/corr/abs-1810-12894, DBLP:journals/corr/BellemareSOSSM16]. To address this, intrinsic rewards are commonly used to promote exploration.
@@ 3 Methods
3.1 Adaptive Entropy Regularization: Achieving a Balance between Exploration and Exploitation
@@
3.2 Motion Continuity Regularization for Quadrupeds: Ensuring Smoothness and Stability
The motion of a robot is continuous. To ensure system stability its dynamic behavior must satisfy constraints [6696874]. In our experiments, we observed that action sequences containing sharp transitions or high-frequency components frequently led to loss of balance and unstable locomotion.
@@ 3.2 Motion Continuity Regularization for Quadrupeds: Ensuring Smoothness and Stability
3.3 Intrinsic Rewards for Improved Exploration–Exploitation and Policy Diversity
@@ 4.1 Experimental Setup
We adapt the original PPO configuration of Rudin et al. for ANYmal in Isaac Gym as our baseline. For comparison, we use the benchmark source PPO with the same architecture, observation and action spaces. For controlled reasons, we compare the series of benchmarks based on the same pipeline, and our model differs only by the addition of adaptive entropy control, motion continuity regularization and intrinsic motivation components. Under this design, any systematic difference in performance between PPO and ECIM can be attributed to the proposed pipeline rather than differences in architecture, training budget or simulator settings.
@@ 4.2 Evaluation Metrics
4.3 ECIM vs PPO: Entropy-Controlled Intrinsic Motivation in Complex Terrains
@@
4.3.1 Learning Efficiency and Task Reward under Entropy-Controlled Exploration
@@
4.3.2 Locomotion Stability and Smoothness in Intrinsically Motivated Gaits
@@
4.3.3 Energy Efficiency of Entropy-Controlled Quadruped Policies
@@
4.4 Dissecting ECIM: Ablations of Entropy Control, Smoothness, and Intrinsic Motivation
@@
4.4.1 Baseline Comparison: PPO vs Full Entropy-Controlled Intrinsic Motivation
@@
4.4.2 Removing Adaptive Entropy Control (AECPOM): Effects on Exploration and Generalization
@@
4.4.3 Removing Motion Continuity (MCRF): Effects on Locomotion Stability and Smoothness
@@
4.4.4 Removing Intrinsic Motivation–Driven Exploration (IMDEEM): Effects on Energy Use
@@
4.4.5 Model Attribution and Cross-Terrain Generalization Robustness
@@ 5 Conclusion
Author Contributions Wanru Gong and Xinyi Zheng conceived and designed the study. Wanru Gong and Xinyi Zheng conducted data gathering. GH performed statistical analyses. Wanru Gong, Xinyi Zheng,Yuan Hui, Xiaopeng Yang and Xiaoqing Zhu wrote the article.